# Time-Optimal Flight with Safety Constraints and Data-driven Dynamics

### v1

@@ III-B Model Predictive Contouring Control
Consider the discrete-time dynamic system of a quadrotor with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. Let us denote the time discretized evolution of the system $\bm{\hat{f}}:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that
$\bm{x}_{k+1}=\bm{\hat{f}}(\bm{x}_{k},\bm{u}_{k}),$
where the sub-index $k$ is used to denote states and inputs at time $t_{k}$.
The general Optimal Control Problem considers the task of finding a control policy $\pi(\bm{x})$, a map from the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized.
@@
$\bm{x}_{k}\in\mathcal{X},\quad\bm{u}_{k}\in\mathcal{U}.$ (3)
In MPCC, the goal is to compromise between maximizing progress along a predefined path, while tracking it accurately. The main ingredients of the cost function are a progress term $\theta$, the contour error $\bm{e^{c}}(\theta)$, and a lag error $\bm{e^{l}}(\theta)$, which describe the perpendicular and tangential error between the current position and its projection on the reference path
$J_{MPCC}(\bm{x})=\sum_{k=0}^{N}\|\bm{e^{l}}(\theta_{k})\|_{Q_{l}}^{2}+\|\bm{e^{c}}(\theta_{k})\|_{Q_{c}}^{2}-\mu\theta_{k}.$ (4)
The resulting optimal control problem is formulated as follows:
@@
$\bm{x}_{N}\in\bm{p}^{d}(\theta_{N}),$ (8)
where $v_{\theta}$ is the first derivative of $\theta$ with respect to time, and $\bm{p}^{d}(\theta_{N})$ is the terminal set which computation will be detailed in the next section. Note that we model the progress $\theta$ as a first order system in order to penalize the variation in progress in the cost, which provides a much smoother state. The norms of the form $\|\cdot\|_{A}^{2}$ represent the weighted Euclidean inner product $\|\bm{v}\|^{2}_{A}=\bm{v}^{T}A\bm{v}$. For more details about this formulation we refer the reader to [14].
@@ IV Methodology
In this section we introduce our enhancements of the standard MPCC formulation introduced in Section III-B.
IV-A Safety set
In this section, we formally define the safety set as a spatial constraint which consistently prevents gate collisions. We then define a periodic terminal set and show that the proposed controller is inherently robust.
We define the safety set as a prismatic tunnel that joins the inner corners of the gates (see Fig. 2). We refer to this formulation as Safe MPCC (SMPCC) to distinguish it from the baseline MPCC introduced in [14]. Two components are required to define such a set: i) a centerline which passes through the gate centers and matches the first derivative with the gate normal; ii) a parameterization of the width and height of the cross section as a function of the progress $\theta$, i.e. $W(\theta_{k})$, $H(\theta_{k})$ respectively.
We employ a simple hermetian spline as the centerline. The gate cross section is then patched around the centerline using standard Frenet-Serret formulas, similar to [20]. At every point of the curve, consider the reference frame that consists of the vectors $[\bm{t}(\theta_{k}),\bm{n}(\theta_{k}),\bm{b}(\theta_{k})]$. Let $\bm{p}_{k}$ be the coordinates of the platform at current time $k$, and let $\bm{p}^{d}(\theta_{k})$ be the centerline. Given the width $W(\theta_{k})$ and height $H(\theta_{k})$ of the tunnel at that point (as shown in Fig. 2), and given the bottom left corner of the tunnel $\bm{p}_{0}(\theta_{k})=\bm{p}^{d}(\theta_{k})-W(\theta_{k})\cdot\bm{n}(\theta_{k})-H(\theta_{k})\cdot\bm{b}(\theta_{k})$, we define the four linear constraints as:
@@ IV-A Safety set
The constraints in (9) are added to problem (8), and are symbolically represented in Fig. 2. Furthermore, we parameterize the width and height of the tunnel by a nominal value $W_{n}$ and an inner gate value $W_{gate}$. A sigmoid function is employed to smoothly transition between these two levels.
This terminal set is has been chosen to be a feasible trajectory that goes through the center of all gates:
$\bm{x}_{N}\in\bm{p}^{d}(\theta_{N}).$ (10)
Figure 2: Safety constraints in position in the shape of a tunnel. The size of the tunnel are set exactly the same as the dimensions of the interior of the gates where at the gates and larger everywhere else.
Such a set terminal set can be for example computed by solving the following optimization problem
@@
$\bm{x}_{k}\in\mathcal{X},\quad\bm{u}_{k}\in\mathcal{U},$
where $M$ is the number of points that parameterize the terminal set, and the first constraints requires the trajectory to be periodic. The trajectory $\bm{p}^{d}(\theta_{k})$ is guaranteed to be positive invariant. Thanks to this terminal set, it is possible to state the following proposition.
@@
The MPCC formulation in (8) is recursively feasible and satisfies constraints at all times.
The proof follows from standard MPC recursive feasibility arguments [59]. It is possible to show that controller (8) is also inherently robust to disturbances by following the same arguments used in [60].
@@ IV-B Dynamics augmentation
In this section we explain how to augment the nominal dynamics $\bm{f}(\bm{x},\bm{u})$ introduced in Section III-A, with a residual term $\bm{g}(\bm{x},\bm{u})$ that captures unmodeled effects.
@@
We first provide a general overview of the forces and torques that act on the system, then identify the sources of model mismatch, and propose a set of polynomial features to approximate these terms from real world data.
@@
We differentiate between the lift force $\bm{f}_{prop}$ produced by the propellers, and the collective aerodynamic forces $\bm{f}_{aero}$ which encompass effects such as drag, induced lift, and blade flapping. The total torque acting on the system is composed of four elements: the torque generated by the individual propellers $\bm{\tau}_{prop}$, the yaw torque $\bm{\tau}_{mot}$ arising from changes in motor speeds, the aerodynamic torque $\bm{\tau}_{aero}$ that captures a variety of aerodynamic influences, and an inertial term $\bm{\tau}_{iner}$.
@@
While $\bm{f}_{prop}$ and $\bm{\tau}_{prop}$ can be precisely estimated from first principles, accurately modelling the aerodynamic forces is significantly more challenging. In [32], the forces acting on the individual propellers were modelled analytically using Blade Element Momentum Theory (BEM), which, due to its precision, notably raises computational demands. Instead, we adopt the methodology from [6] in which simple, low-dimensional polynomial features are fitted to real world data through conventional regression techniques. This method operates on the premise that aerodynamic effects are primarily characterized by the drone’s linear body velocity, and the average of squared (four) individual motor speeds. The features are then selected by combining the individual terms $v_{x}$, $v_{y}$, $v_{z}$, $\Omega^{2}$ in different powers:
$f_{x}$ $\sim v_{x}+v_{x}^{3}+\Omega^{2}+v_{x}\Omega^{2}$
$f_{y}$ $\sim v_{y}+v_{y}^{3}+\Omega^{2}+v_{y}\Omega^{2}$
$f_{z}$ $\sim v_{z}+v_{z}^{3}+v_{xy}+v_{xy}^{2}+v_{xy}\Omega^{2}+v_{z}\Omega^{2}+v_{xy}v_{x}\Omega^{2}$
$\tau_{x}$ $\sim v_{y}+\Omega^{2}+v_{y}\Omega^{2}$
$\tau_{y}$ $\sim v_{x}+\Omega^{2}+v_{x}\Omega^{2}$
$\tau_{z}$ $\sim v_{x}+v_{y}.$ (11)
The respective coefficients are identified using force and torque measurements that are directly derived onboard IMU motorspeed and IMU measurements as well as from ground truth data, captured with a VICON [^1] https://www.vicon.com/ system. The data is gathered from the same race track to ensure that it captures the unique aerodynamic effects that arise from the racing maneuvers. It’s important to note that the coefficients are identified offline and remain constant at runtime. Additionally, the residual model requires an accurate estimate of the motor speeds, which are not provided in the original MPCC formulation from Problem 8. To address this, we modfy the initial problem to include a first-order motor model within the nominal dynamics as outlined in Eq. (1)
$\dot{\bm{\Omega}}=\frac{1}{\tau_{mot}}(\bm{\Omega_{des}}-\bm{\Omega}),$ (12)
where $\bm{\Omega_{des}}$ and $\bm{\Omega}$ denote the desired and the actual motor speeds, respectively.
@@ IV-C TuRBO tuning
We consider the SMPCC controller as a black box function in the context of Bayesian Optimization. For a given set of controller parameters $\bm{\phi}$, we run an episode, collect a trajectory $\tau(\bm{\phi})$, and compute the reward $R(\tau(\bm{\phi}))$. For the sake of clarity, we will omit the trajectory from the notation and simply refer to the reward as $R(\bm{\phi})$. The controller tuning task can be framed as an optimization problem:
$\bm{\phi}_{n+1}=\arg\max_{\bm{\phi}\in\Theta}R(\bm{\phi}).$ (13)
Problem (13) presents two caveats: i) no analytical form of $R(\bm{\phi})$ exists, allowing only for pointwise evaluation; ii) each evaluation corresponds to completing an entire episode within the simulator, which means the number of evaluations is capped by our interaction budget with the simulated environment. BO works in an iterative manner, relying on two key ingredients: i) a probabilistic surrogate model approximating the objective function and ii) an acquisition function $\alpha(\bm{\phi})$ that determines new evaluation points, balancing exploration and exploitation. A common choice for the surrogate model are Gaussian Processes (GPs), which allow for closed-form inference of the posterior mean $\mu(\bm{\phi})$ and variance $\sigma^{2}(\bm{\phi})$. The next evaluation point $\bm{\phi}_{n+1}$ is then chosen by maximizing the acquisition function $\alpha(\bm{\phi})$. We utilize the Upper Confidence Bound (UCB) acquisition function, defined as $\alpha_{UCB}(\bm{\phi})=\mu(\bm{\phi})+\beta\sigma(\bm{\phi})$, where $\beta$ is the exploration parameter. This process is iterated until the evaluation budget is exhausted. Among the various BO methodologies, we opt for TuRBO [47], a global BO strategy that runs multiple independent local BO instances concurrently. Each local surrogate model enjoys the benefits of local BO such as diverse modeling of the objective function across different regions. Each local run explores within a Trust Region (TR) - a polytope centered around the optimal solution of the local instance. The base side length of the TR, $L$ is adjusted based on the success rate of evaluations, to guide exploration towards promising areas. In this work we use the TuRBO-1 implementation from the BOTorch library [61] and modify it for an eightfold setup, TuRBO-8.
In contrast to the baseline MPCC [14], which requires $N_{\bm{\phi}}=2n_{gates}+4$ parameters to be tuned due to the complexity of the contour function, the SMPCC reduces the number of tunable parameters to $N_{\bm{\phi}}=8$. From Problem 8, the parameters in question within the SMPCC include $Q_{l}$, $Q_{c}$, $Q_{\omega}$, $R_{v_{\theta}}$, $R_{\Delta f}$, and $\mu$, with $Q_{c}$, $Q_{\omega}$ further divided into horizontal and vertical components.
@@
$R(\bm{\phi})=-\frac{1}{M}\sum_{i=0}^{M}t_{i}-\gamma r_{fail},$ (14)
where $t_{i}$ denotes the lap time of the i-th lap, while $r_{fail}$ represents the solver failure rate, i.e. the proportion of steps that fail relative to the total steps. We apply a scaling factor of $\gamma=100$ to the penalty term. Such failures arise mainly due to infeasibilities when the system is pushed to its constraint bounds. Incorporating this penalty term has proven effective in practice, as it smoothens the objective function in favor of parameters that minimize solver failures. This reward structure is consistently applied across all experiments.
Figure 3: Overview of the complete pipeline. Starting with an initial sample of tuning parameters $\bm{\phi}$, a trajectory is run. TuRBO then receives the reward associated to the collected trajectory and then updates the tuning parameters for the next sample. For the residual model learning, data is collected from the real quadrotor system and $\bm{g}(\bm{x},\bm{u})$ is fitted to that data.
@@ V Experiments
In this section, we compare our proposed method, SMPCC, against two baselines: RL and MPCC, and conduct a series of ablation studies. The methodology is consistent across all experiments, and utilizes the same quadrotor configuration. We choose the Split-S race track, featuring 7 gates, for all our experiments due to its prevalent use in previous works [13, 14, 15, 7]. We test the performance of our approach across three distinct simulation environments: i) a simple simulator which uses the nominal dynamics; ii) a high-fidelity simulator that calculates the propeller forces via Blade Element Momentum Theory (BEM) [32]; and iii) a data-driven simulator that predicts aerodynamic forces from real world data. We then validate our method in the real world. All experiments were executed on the same hardware under uniform conditions.
@@
SMPCC (ours) Nominal TurBO 5.16 $\pm$ 0.02 100 5.30 $\pm$ 0.02 100 5.37 $\pm$ 0.09 100 5.41 $\pm$ 0.14 100
@@
RL - - 5.14 $\pm$ 0.09 100 - - 5.26 $\pm$ 0.32 100 5.35 $\pm$ 0.15 85.0
TABLE I: Results for MPCC, SMPCC and RL. There are four columns depending on which environment the experiment was conducted in. Three of these columns are for different simulation environments, and the last column for real world results. We can see how for our SMPCC, the success rate in all simulations and in the real world is 100%.

### v2

@@ III-B Model Predictive Contouring Control
We consider the discrete-time dynamic system of a quadrotor with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. We denote the time discretized evolution of the system $\bm{\hat{f}}:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that:
$\bm{x}_{k+1}=\bm{\hat{f}}(\bm{x}_{k},\bm{u}_{k})$ (3)
where the index $k$ refers to the states and inputs at time $t_{k}$. The general Optimal Control Problem (OCP) considers the task of finding a control policy $\pi(\bm{x})$, a map from the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized:
@@
$\bm{x}_{k}\in\mathcal{X},\quad\bm{u}_{k}\in\mathcal{U}$ (4)
In MPCC, the goal is to compromise between maximizing progress along a predefined path, while tracking it accurately. The main ingredients of the cost function are the progress term $\theta$, the contour error $\bm{e^{c}}(\theta)$, and the lag error $\bm{e^{l}}(\theta)$, which describe the perpendicular and tangential error between the current position and its projection on the reference path:
$J_{MPCC}(\bm{x})=\sum_{k=0}^{N}\|\bm{e^{l}}(\theta_{k})\|_{Q_{l}}^{2}+\|\bm{e^{c}}(\theta_{k})\|_{Q_{c}}^{2}-\mu{v_{\theta_{k}}}$ (5)
The resulting OCP is formulated as follows:
@@
where $v_{\theta}$ is the first derivative of $\theta$ with respect to time. We model the progress $\theta$ as a first order system to penalize the variation in progress in the cost, which provides a much smoother state. The norms of the form $\|\cdot\|_{A}^{2}$ represent the weighted Euclidean inner product $\|\bm{v}\|^{2}_{A}=\bm{v}^{T}A\bm{v}$. For more details about this formulation we refer the reader to [14].
@@ IV Methodology
In this section we introduce our enhancements to the standard MPCC formulation described in Section III-B.
IV-A Safety constraints
We first introduce the track constraint as a spatial constraint which consistently prevents gate collisions. We then define a terminal set and show that the proposed controller is recursively feasible.
We define the track constraint as a prismatic tunnel that joins the inner corners of the gates (see Figure 1). We refer to this formulation as MPCC++ to distinguish it from the baseline MPCC introduced in [14]. Two components are required to define such a set: i) a centerline which passes through the gate centers and matches the first derivative with the gate normal; and ii) a parameterization of the width and height of the cross section as a function of the progress $\theta$, i.e. $W(\theta_{k})$, $H(\theta_{k})$ respectively.
We employ a simple hermetian spline as the centerline. The gate cross section is constructed around the centerline using standard Frenet-Serret formulas, similar to [19]. At every point of the curve, consider the reference frame that consists of the vectors $[\bm{t}(\theta_{k}),\bm{n}(\theta_{k}),\bm{b}(\theta_{k})]$. Let $\bm{p}_{k}$ be the coordinates of the platform at current time $k$, and $\bm{p}^{d}(\theta_{k})$ the corresponding point on the centerline. For the width $W(\theta_{k})$ and height $H(\theta_{k})$ of the tunnel, define the bottom left corner of the tunnel as $\bm{p}_{0}(\theta_{k})=\bm{p}^{d}(\theta_{k})-W(\theta_{k})\cdot\bm{n}(\theta_{k})-H(\theta_{k})\cdot\bm{b}(\theta_{k})$. The boundary of the tunnel is then determined by four halfspace constraints as follows:
@@ IV-A Safety constraints
Furthermore, we set $W(\theta_{k})=H(\theta_{k})$ and parameterize $W(\theta_{k})$ by two distinct values: a nominal value $W_{n}$ for the broader sections of the tunnel, and an inner gate value $W_{gate}$ for the narrower sections at the gates. A sigmoid function is employed to smoothly transition between these two levels, ensuring a gradual narrowing of the tunnel as it approaches a gate.
Figure 1: Spatial constraint forming a tunnel around the centerline. The width and height are parameterized by $W(\theta_{k})$ and $H(\theta_{k})$, respectively.
We define the terminal set $\mathcal{X}_{f}$ as a periodic feasible trajectory which passes through the center of the gates, similar to [58]:
$\bm{p}_{N}\in\mathcal{X}_{f}$ (11)
where $\bm{p}_{N}$ is the position of the last shooting node $\bm{x}_{N}$. By making the first and last elements of the trajectory coincide, we establish its periodicity, which ensures that the resulting trajectory is a positive invariant set. We compute this set offline by solving the following optimization problem:
@@
$\bm{x}_{k}\in\mathcal{X},\quad\bm{u}_{k}\in\mathcal{U}$ (12)
where $M$ is the number of points that parameterize the terminal set, and the first constraint ensures periodicity of the trajectory. The resulting discrete terminal set is obtained from the solution of Problem 12:
$\mathcal{X}_{f}=\{\bm{p}_{0},\bm{p}_{1},\cdots,\bm{p}_{M}\}$ (13)
where $\bm{p}_{i}$ is the position of the i-th shooting node.
@@
The MPCC formulation in Problem 9 subject to constraints (10) and (11) is recursively feasible and satisfies constraints at all times.
The proof follows from standard MPC recursive feasibility arguments [59]. It is also possible to show that Problem 9 is inherently robust to disturbances by following the same arguments used in [58].
@@ IV-B Dynamics augmentation
In this section we augment the nominal dynamics $\bm{f}(\bm{x},\bm{u})$ introduced in Section III-A, with a residual term $\bm{g}(\bm{x},\bm{u})$ that captures unmodeled effects. We first provide a general overview of the forces and torques that act on the system, then identify the sources of model mismatch, and propose a set of polynomial features to approximate these terms from real-world data.
@@
We differentiate between the lift force $\bm{f}_{prop}$ produced by the propellers, and the collective aerodynamic forces $\bm{f}_{aero}$ which encompasses effects such as drag, induced lift, and blade flapping. The total torque acting on the system is composed of four elements: i) the torque generated by the individual propellers $\bm{\tau}_{prop}$; ii) the yaw torque $\bm{\tau}_{mot}$ arising from changes in motor speeds; iii) the aerodynamic torque $\bm{\tau}_{aero}$ that captures a variety of aerodynamic influences; and iv) the inertial term $\bm{\tau}_{iner}$.
@@
While $\bm{f}_{prop}$ and $\bm{\tau}_{prop}$ can be precisely estimated from first principles, modelling the aerodynamic terms is significantly more challenging. In [34], $\bm{f}_{prop}$ and $\bm{\tau}_{prop}$ were modelled analytically using Blade Element Momentum Theory (BEM), which due to its precision, notably raises computational demands. Instead, we follow the structure proposed by [6], in which a simple polynomial model is fitted to real-world data through conventional regression techniques. The polynomial model describes the aerodynamic forces and torques as a linear combination of polynomial features involving the drone’s linear velocity (in body-frame) and the mean-squared rotor speed, $\Omega^{2}$.
$f_{x}$ $=\bm{C}_{f_{x}}\begin{bmatrix}v_{x}&v_{x}^{3}&\Omega^{2}&v_{x}\Omega^{2}\end{bmatrix}^{T}$
$f_{y}$ $=\bm{C}_{f_{y}}\begin{bmatrix}v_{y}&v_{y}^{3}&\Omega^{2}&v_{y}\Omega^{2}\end{bmatrix}^{T}$
$f_{z}$ $=\bm{C}_{f_{z}}\begin{bmatrix}v_{z}&v_{z}^{3}&v_{xy}&v_{xy}^{2}&v_{xy}\Omega^{2}&v_{z}\Omega^{2}&v_{xy}v_{z}\Omega^{2}\end{bmatrix}^{T}$
$\tau_{x}$ $=\bm{C}_{\tau_{x}}\begin{bmatrix}v_{y}&\Omega^{2}&v_{y}\Omega^{2}\end{bmatrix}^{T}$
$\tau_{y}$ $=\bm{C}_{\tau_{y}}\begin{bmatrix}v_{x}&\Omega^{2}&v_{x}\Omega^{2}\end{bmatrix}^{T}$
$\tau_{z}$ $=\bm{C}_{\tau_{z}}\begin{bmatrix}v_{x}&v_{y}\end{bmatrix}^{T}$ (14)
where $v_{xy}$ represents the horizontal velocity component. The selection of these terms is grounded in their relevance to the Blade Element Momentum theory (BEM) solution [34, 6], reflecting their significant impact on the system dynamics while maintaining computational efficiency. The ground truth force and torque values are derived from IMU measurements, as well as from a VICON [^1] https://www.vicon.com system. The respective coefficients $\bm{C}_{f}$, $\bm{C}_{\tau}$ are then identified using ordinary linear least-squares regression, and remain constant at runtime. The data is gathered from the same race track to ensure that it captures the unique aerodynamic effects that arise from the racing maneuvers.
@@ IV-C TuRBO tuning
We consider the MPCC++ controller as a black-box function in the context of BO. For a given set of controller parameters $\bm{\phi}$, we run an episode, collect a trajectory $\tau(\bm{\phi})$, and compute the reward $R(\tau(\bm{\phi}))$. For the sake of clarity, we simply refer to the reward as $R(\bm{\phi})$. The controller tuning task can be framed as an optimization problem:
$\bm{\phi}_{n+1}=\arg\max_{\bm{\phi}\in\Theta}R(\bm{\phi})$ (15)
Problem 15 presents two caveats: i) no analytical form of $R(\bm{\phi})$ exists, allowing only for pointwise evaluation; and ii) each evaluation corresponds to completing an entire episode within the simulator, which means the number of evaluations is capped by our interaction budget with the simulated environment. BO works in an iterative manner, relying on two key ingredients: i) a probabilistic surrogate model approximating the objective function; and ii) an acquisition function $\alpha(\bm{\phi})$ that determines new evaluation points, balancing exploration and exploitation. A common choice for the surrogate model are Gaussian Processes (GPs), which allow for closed-form inference of the posterior mean $\mu(\bm{\phi})$ and variance $\sigma^{2}(\bm{\phi})$. The next evaluation point $\bm{\phi}_{n+1}$ is then chosen by maximizing the acquisition function $\alpha(\bm{\phi})$. We utilize the Upper Confidence Bound (UCB) acquisition function, defined as $\alpha_{UCB}(\bm{\phi})=\mu(\bm{\phi})+\beta\sigma(\bm{\phi})$, where $\beta$ is the exploration parameter. This process is iterated until the evaluation budget is exhausted. Among the various BO methodologies, we opt for TuRBO [60], a global BO strategy that runs multiple independent local BO instances concurrently. Each local surrogate model enjoys the benefits of local BO such as diverse modeling of the objective function across different regions. Each local run explores within a Trust Region (TR) - a polytope centered around the optimal solution of the local instance. The base side length of the TR, $L$, is adjusted based on the success rate of evaluations to guide exploration towards promising areas. In this work we use the TuRBO-1 implementation from the BOTorch library [61] and modify it for an eightfold setup, TuRBO-8.
In contrast to the baseline MPCC [14], which requires $N_{\bm{\phi}}=2n_{gates}+4$ parameters to be tuned due to the complexity of the contour function, MPCC++ reduces the number of tunable parameters to $N_{\bm{\phi}}=8$. From Problem 9, the parameters in question within the MPCC++ formulation include $Q_{l}$, $Q_{c}$, $Q_{\omega}$, $R_{v_{\theta}}$, $R_{\Delta f}$ and $\mu$, with $Q_{c}$, $Q_{\omega}$ further divided into horizontal and vertical components.
@@
$R(\bm{\phi})=-\frac{1}{M}\sum_{i=0}^{M}t_{i}-\gamma r_{fail}$ (16)
where $t_{i}$ denotes the lap time of the i-th lap, while $r_{fail}$ represents the solver failure rate, i.e. the proportion of steps that fail relative to the total steps. We apply a scaling factor of $\gamma=100$ to the penalty term. Such failures arise mainly due to i) the complexity of the optimization problem at hand; and ii) the need to solve the optimization problem in real-time (one SQP iteration). Incorporating this penalty term, while not strictly necessary, has proven effective in practice, as it smoothens the objective function in favor of parameters that minimize solver failures. This reward structure is consistently applied across all experiments.
@@ V Experiments
In this section, we compare our proposed method, MPCC++, against two baselines: MPCC and RL, and conduct a series of variations in the formulation. The methodology is consistent across all experiments, and utilizes the same quadrotor configuration. We choose the Split-S race track, featuring 7 gates, for all our experiments due to its prevalent use in previous works [7, 13, 14, 15]. We test our approach across three distinct simulation environments: i) a simple simulator which uses the nominal dynamics; ii) a high-fidelity simulator that calculates the propeller forces via Blade Element Momentum Theory (BEM) [34]; and iii) a data-driven simulator that predicts aerodynamic forces from real-world data. We then validate our method in the real world. All experiments were conducted on the same hardware under uniform conditions.
@@
MPCC++ (ours) Nominal TuRBO 5.16 $\pm$ 0.02 100 5.30 $\pm$ 0.02 100 5.37 $\pm$ 0.09 100 5.41 $\pm$ 0.14 100
@@
RL [7] - - 5.14 $\pm$ 0.09 100 - - 5.26 $\pm$ 0.32 100 5.35 $\pm$ 0.15 85.0
TABLE I: Results for MPCC, MPCC++, and RL across three simulation environments - Simple, BEM, and Residual - as well as for real-world experiments. MPCC++ achieves a 100% success rate across all simulation and real-world environments.